The marginal revenue of a company is given by MR-40 +60 +


602, where Q is the number of units sold for a period. Find the total revenue


function if at Q-0 total revenue is equal to 0.


1
Expert's answer
2021-07-12T15:54:09-0400

"TR=\int MRdQ=\int^Q_0(40Q^2+60Q+602)dQ=40Q^3/3+30Q^2+602Q"